OpenStudy (zuha):

how can I convert 2.96 in to a fraction

OpenStudy (anonymous):

\[2.96 = \frac{296}{100}\] You can simplify that of course.

OpenStudy (zuha):

how?

OpenStudy (zuha):

help!!

OpenStudy (anonymous):

\[\frac{296}{100} = \frac{\cancel{2} \times 148}{\cancel{2} \times 50}\] Keep going...

OpenStudy (zuha):

I came to the answer 37/100

OpenStudy (anonymous):

You have to divide the denominator, too.

OpenStudy (texaschic101):

296/100 = 148/50 = 74/25

OpenStudy (zuha):

So it's mixed number is 2 24/25, right?

OpenStudy (texaschic101):

yes

OpenStudy (zuha):

have another Q Im attaching it

OpenStudy (zuha):

please guide me

OpenStudy (texaschic101):

what do you need to do ?

OpenStudy (zuha):

match each of the point with number of the fraction

OpenStudy (zuha):

please check the attachment

OpenStudy (anonymous):

The attachment just says 5/6 ...??

OpenStudy (texaschic101):

the attachment just says 5/6

OpenStudy (zuha):

im drawing it

OpenStudy (zuha):



OpenStudy (zuha):

Sorry i messed up the options in copying ! here they are A=five over six, B=five over sixteen, C=seven over twelve, D=one over eight A=one over eight, B=five over sixteen, C=seven over twelve, D=five over six A=seven over twelve, B=five over six, C=five over sixteen, D=one over eight A=five over sixteen, B=one over eight, C=seven over twelve, D=five over six

OpenStudy (texaschic101):

5/6 is 1 1/6 I am having trouble with your line graph....what is the middle line supposed to be ? What number ?

OpenStudy (zuha):

Please help me

OpenStudy (zuha):

It is soupposed to be the middle of the line

OpenStudy (texaschic101):

I suggest turning the fractions to decimals, then you can see which one is smallest and which one is biggest. 5/6 = 0.83 5/16 = 0.31 7/12 = .58 1/8 = 0.125 A. = 1/8 B = 5/16 C = 7/12 D = 5/6

OpenStudy (zuha):

good thats what I thought Thanks

OpenStudy (texaschic101):

anytime :)